A zero-coupon bond with face value $1,000 and maturity of four years sells for $747.22.

a. What is its yield to maturity? (Round your answer to 2 decimal places.)

Yield to maturity             %

b. What will the yield to maturity be if the price falls to $731? (Round your answer to 2 decimal places.)

Yield to maturity             %
Answer #1

a. Yield to Maturity :

Present Value = Future Value / ( 1+ YTM)^ n

747.22 = 1000 /(1+YTM)^4

[(1000/747.22) ^(1/4) ] - 1 =YTM

YTM = 7.56%

Answer : 7.56%

b.

Yield to Maturity :

Present Value = Future Value / ( 1+ YTM)^ n

731 = 1000 /(1+YTM)^4

[(1000/731) ^(1/4) ] - 1 =YTM

YTM = 8.15%

Answer : 8.15%